Calculate the velocity of water in a pipe of cross sectional area 0.02sqaure meters which delivers 115 liters of water per second.
1
Expert's answer
2019-12-11T10:36:42-0500

The flux of water having velocity "v" through pipe of cross sectional area "S" in time "t" is "I = S v t".

Hence, the velocity is "v = \frac{I}{S t} = \frac{115 \cdot 10^{-3} m^3}{0.02 m^2 \cdot 1 s} = 5.75 \frac{m}{s}".